# Satellite Drag Analysis During the May 2024 Geomagnetic Storm

### v1

@@ 1 Introduction
Figure 1 shows a worrying trend in space operations. First, the number of payloads launched into Earth’s orbit has grown dramatically since the last major geomagnetic storm (the number of active payloads is somewhat contested, so the annual number of payloads launched, recorded in [12], is a better proxy for the change in activity). This growth is largely attributable to proliferated LEO constellations like Starlink and OneWeb, but also to the transition towards inexpensive small satellites capable of being launched as ride-shares to orbit. Such a large increase in traffic throughout LEO, along with a ballooning debris population from previous fragmentation events (especially the notable events in 2008 [13], 2009 [14], and 2021 [15]) has made satellite conjunction assessment and collision avoidance a necessary capability on most new spacecraft to ensure satellite safety.
Geomagnetic storms are most common near solar maximum, the period of peak activity in the roughly 11-year solar cycle. Sun spots, solar flares, and CMEs happen more frequently during solar maximum. Solar extreme ultraviolet (EUV) radiation, the main driver for heating and ionization in the upper atmosphere, also peaks at this time. During major storms in 1989 and 2003, NORAD lost track of many satellites for several days [16]. A similar failure today might have dire consequences. Figure 1 shows that while there is precedent for geomagnetic storms from 1989 and 2003, the May 2024 storm is unique in that it is the first to occur during a new paradigm in satellite operations in LEO. As the solar cycle continues to peak throughout 2024 and 2025, continued disruptions to operations are likely to occur.
@@ 4 Operational Implications
For many operators managing satellites during the storm, such a sudden drop in orbital altitude is untenable. Unplanned orbital decay can disrupt constellations by causing uneven satellite altitudes, which results in undesirable orbit phasing in the short term and relative plane drift over the long term. Other satellites performing Earth observation tasks may also have similarly tight constraints on orbital altitude and require regular station-keeping. Figure 6 shows the number of LEO tracked objects maneuvering over time with the time history of in the background for reference. The figure includes both the May 2023 storm and the October 2003 Halloween storm. A maneuver is counted when the time-averaged altitude of an object increases by > 100 m over a 3-hour window. In the May 2023 storm, about 1000 of the nearly 10,000 active payloads in LEO appear to be maneuvering during the quiet period leading up to the storm. After the storm hits, with some offset to account for the time it takes for drag decay to accumulate, thousands of satellites begin to maneuver en masse in response to the sudden increase in atmospheric density. For comparison, there was no discernable change in maneuver activity in LEO during the October 2003 Halloween storm. Most of the May 2024 maneuver activity is attributable to the Starlink constellation, which performs autonomous orbit maintenance and thus responds quickly to perturbing events. Onboard orbit maintenance will become more common as other proliferated LEO constellations are established.
@@ 6 Conclusion
This note highlighted impacts from the May 2024 geomagnetic storm on satellite operations in LEO. The storm represented a serious challenge for the existing conjunction assessment infrastructure as it produced large, unpredictable perturbations on satellite trajectories in LEO. New proliferated LEO constellations require tight station-keeping bounds to prevent undesirable orbit phasing. Automated station-keeping, especially from the Starlink constellation, caused nearly half of all the active satellites in LEO to maneuver at once in response to the storm. The combination of unpredictable satellite drag and bulk maneuvering made it very difficult or impossible to identify potential conjunctions during the storm and in the days that followed. While the storm represented a risk to the LEO environment in the short term, it also helped to hasten the removal of debris populations from orbit. This passive debris removal is critical for the long-term sustainability of operations in LEO. Moving forward, it is important that we recognize the limits that the environment imposes on satellite activity in LEO. Operators and regulators should consider the robustness of the conjunction assessment infrastructure to events of this kind when deciding how much to rely on it.
@@ 7 Acknowledgments
This work was supported by the National Science Foundation Graduate Research Fellowship under Grant No. 1745302. The authors gratefully acknowledge the sponsor for their support.

### v2

# Satellite Drag Analysis During the May 2024 Gannon Geomagnetic Storm

@@ 1 Introduction
Figure 1 shows a worrying trend in space operations. First, the number of payloads launched into Earth’s orbit has grown dramatically since the last major geomagnetic storm (the number of active payloads is somewhat contested, so the annual number of payloads launched, recorded at space-track.org, is a better proxy for the change in activity). This growth is largely attributable to proliferated LEO constellations like Starlink and OneWeb, but also to the transition towards inexpensive small satellites capable of being launched as ride-shares to orbit. Such a large increase in traffic throughout LEO, along with a ballooning debris population from previous fragmentation events (especially the notable events in 2008 [12], 2009 [13], and 2021 [14]) has made satellite conjunction assessment and collision avoidance a necessary capability on most new spacecraft to ensure satellite safety.
Geomagnetic storms are most common near solar maximum, the period of peak activity in the roughly 11-year solar cycle. Sun spots, solar flares, and CMEs happen more frequently during solar maximum. Solar extreme ultraviolet (EUV) radiation, the main driver for heating and ionization in the upper atmosphere, also peaks at this time. During major storms in 1989 and 2003, NORAD lost track of many satellites for several days [15]. A similar failure today might have dire consequences. Figure 1 shows that while there is precedent for geomagnetic storms from 1989 and 2003, the May 2024 Gannon storm is unique in that it is the first to occur during a new paradigm in satellite operations in LEO. As the solar cycle continues to peak throughout 2024 and 2025, continued disruptions to operations are likely to occur.
@@ 4 Operational Implications
For many operators managing satellites during the storm, such a sudden drop in orbital altitude is untenable. Unplanned orbital decay can disrupt constellations by causing uneven satellite altitudes, which results in undesirable orbit phasing in the short term and relative plane drift over the long term. Other satellites performing Earth observation tasks may also have similarly tight constraints on orbital altitude and require regular station-keeping. Figure 6 shows the number of LEO tracked objects maneuvering over time with the time history of in the background for reference. The figure includes both the May 2023 storm and the October 2003 Halloween storm. A maneuver is counted when the mean orbit altitude of an object increases by > 100 m over a 3-hour window. For the May 2023 storm, about 300 of the nearly 10,000 active payloads in LEO appear to be maneuvering during the quiet period leading up to the storm. After the storm hits, with some offset to account for the time it takes for drag decay to accumulate, thousands of satellites begin to maneuver en masse in response to the sudden increase in atmospheric density. For comparison, maneuver activity in LEO was not comparably impacted during the October 2003 Halloween storm. Most of the May 2024 maneuver activity is attributable to the Starlink constellation, which performs autonomous orbit maintenance and thus responds quickly to perturbing events. Onboard orbit maintenance will become more common as other proliferated LEO constellations are established.
@@ 6 Conclusion
This note highlighted impacts from the May 2024 Gannon geomagnetic storm on satellite operations in LEO. The storm represented a serious challenge for the existing conjunction assessment infrastructure as it produced large, unpredictable perturbations on satellite trajectories in LEO. New proliferated LEO constellations require tight station-keeping bounds to prevent undesirable orbit phasing. Automated station-keeping, especially from the Starlink constellation, caused nearly half of all the active satellites in LEO to maneuver at once in response to the storm. The combination of unpredictable satellite drag and bulk maneuvering made it very difficult or impossible to identify potential conjunctions during the storm and in the days that followed. While the storm represented a risk to the LEO environment in the short term, it also helped to hasten the removal of debris populations from orbit. This passive debris removal is critical for the long-term sustainability of operations in LEO. Moving forward, it is important that we recognize the limits that the environment imposes on satellite activity in LEO. Operators and regulators should consider the robustness of the conjunction assessment infrastructure to events of this kind when deciding how much to rely on it.
@@ 7 Acknowledgments
This work was supported by the National Science Foundation Graduate Research Fellowship under Grant No. 1745302 and the joint NSF-NASA Space Weather with Quantified Uncertainties program under NSF grant number 2028125. The authors gratefully acknowledge the sponsors for their support.